UT 20150510 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson (remote)
Observer AO: DMiller (on skype), JChristou (LBTO)
Telescope Operator: SAllenson
LBTO Support: TSargent

Plan for the night

- Photo ZP's

- AO-com

Summary

  • Active flexure compensation works in imaging mode!

  • Photo ZPs now in all filters for N30 camera

  • Lots of TCS issues

  • Aladin issue (again) independent of machine used (remote or luci)

Details

All times are UT
Files are luci2.20150510.NNNN.fits

02:22 Sunset, photometric, T = -2deg, Hum = 50%, wind 5-10m/s, seems to be a good night waiting for us

02:45 Steve checked pointing and collimation, seeing looks good (about 0.7" on guider) JChristou checks AO, yes, works

  • Photometric ZPs

03:02 BS 9142, photo ZPs zJHKKs, N3.75, #0001-

03:44 BS 9142, photo ZP's zJHKKs, N30, #xxx

  • Image quality

04:20 NGC 4147, N30, Ima_Qual, preset failed, TCS is hanging (IT 5559 again), redid preset
failed, GS seems to be a galaxy, need to rewrite script

  • Active flexure test - imaging

04:37 BS9148, preset committed, but not sent to telescope, redid, works,
seeing on guider 0.9", K-band FWHM = 0.09"! SHIT, shell ripped three times,
came from three actuators: 562, 336, 423, recollimated, JChristou tries to close the loop
finally done, continue script, 5*5min with and 5*5min without AFC in a dice5 pattern

Active flexure compensation in imaging mode works

  • PSF- test, AFC-test cont'd

06:50 3C294, difference between mag AO-ref star in catalog and reported by FLAO is about 1mag (FLAO reports fainter)

07:15 All at the sudden, Aladin doesn't want to display recent images (GUIs do NOT freeze), problem
shows also on Alex notebook, restart LUCI-SW

07:53 Redo 3C 294, loop closed, loop opens, closed, open...hotspot has moved, restart GCS etc.pp.
still do not get nice images. Reason most likely: the AO refstar is a double with about 0.15" separation!

09:05 B1422+22, loop closed, images look nice...mistake in script (relative instead of absolute offsets), Dave modified that
and restartet,...TCS hungs again....script restartet...run AO failed. 2 actuators jumped...Steve resets shell
...and restarted...finally started #0083- , Bin=4, 100Hz, 36 modes, AO-Refstar has 14.8mag

10:35 go to zenith to flatten shell

10:40 BS 9177 do the rest of missing ZPs for 4 nB-filters with N30-camera,

Good seeing all night (< 1" on guider)

11:24 Steve is closing up

End of Com#3

-- %USERSIG{JochenHeidt - 2015-05-010}%
Topic revision: r6 - 19 Mar 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback